Z~N (0,1^2) find a) P (Z>2.0) b) P (Z <-2.0) c) P ( -1 <Z <0.5)
1
Expert's answer
2020-06-08T20:17:52-0400

"Z\sim N(0,1/2)" , "\mu =0,\ \sigma ^2=1/2"

We will use the following properties:

"P(X<a)=F(a)=1-F(-a)"

"P(X>a)=1-F(a)=F(-a)"

"P(a<X<b)=F(b)-F(a)"

"F(x)=\Phi (\frac{x-\mu}{\sigma})"


a) "P(Z>2)=1-F(2)=1-\Phi(2\sqrt2)\approx 1-\Phi(2.82)=1-0.9976=0.0024"


b) "P(Z<-2)=F(-2)=1-F(2)=0.0024"


c) "P(-1<Z<0.5)=F(0.5)-F(-1)=\Phi(1/\sqrt 2)-\Phi(-\sqrt 2)\approx\Phi(0.70)- \Phi(-1.41)=0.7580-0.0793\approx 0.68"